# Modeling the Dynamics of Sub-Millisecond Electroadhesive Engagement and Release Times

### v1

@@ IV-D Release Time Results and Discussion
V Simulating Parameter Sweeps for Engagement and Release Time
@@
Fig. 12 shows our model’s predicted engagement time over different drive voltages, substrate widths and lengths, dielectric thicknesses, and drive circuit rise times. As shown, the substrate width, drive circuit rise time, and voltage all have very large effects on engagement time, while the dielectric thickness and substrate length have much smaller effects. The substrate’s width affects engagement dynamics much more than its length does because the damping force increases as the cube of width but linearly with length. We also swept a variety of other parameters that did not have a large impact on the predicted engagement time, largely because our actuators engage so quickly that electrical and mechanical dynamics all change on a similar time scale. Our model predicted that the dielectric constant $\kappa$ would have minimal effect on engagement time, because when the voltage first rises (i.e., when $T_{air}$ is the largest) the $\kappa^{2}$ terms in the numerator and denominator of Eq. 4 roughly cancel out. The preload force also had little effect in our predicted and experimental results, which is promising for future work because we found a small preload significantly reduces wrinkling in the dielectric film.
Figure 12: Our model’s predicted engagement time while varying voltage together with (a) substrate width, (b) dielectric thickness, (c) substrate length, and (d) the 10% - 90% rise time of our high voltage drive circuit. The black dotted lines indicate the values tested in this paper. Unless otherwise specified, the substrate width tested was 2 mm, the drive signal was a 300 V, 1 kHz bipolar square wave, and $F_{preload}=$ 0.125 N.
@@
Figure 13: Our model’s predicted release time while varying voltage together with (a) substrate width, (b) dielectric thickness, (c) substrate length, and (d) the 90% - 10% fall time of our high voltage drive circuit. The black dotted lines indicate the values tested in this paper. Unless otherwise specified, the substrate width tested was 2 mm, the drive signal was a 300 V, 1 kHz bipolar square wave, $F_{preload}=$ 0.125 N, and $F_{shear}(t_{r})/F_{shear,max}$ = 0.8.
@@
Fig. 13 shows our model’s predicted release time over the same design parameters. The substrate width matters the most for release time, followed by voltage and dielectric thickness. For larger substrate widths, after the rate of increase in our contact force $F_{k}(T_{air})$ at small air gaps outpaces that of our EA force $F_{ea}(T_{air})$, larger substrate widths are predicted by our model to see lower release times. The dielectric thickness has an interesting local minima in its graph as the dielectric’s increasing mass compensates for the loss of EA force capacity when determining the initial air gap before release. Diller et al. [6] similarly observed higher release times at larger substrate widths and lengths, lower voltages, and in an intermediate dielectric thickness range for parallel plate EA clutches, but to our knowledge this is the first time these trends have been replicated by a physics-based model. The high voltage drive circuitry’s fall time only has a small positive correlation with release time, as it slightly delays when the load cell can overcome static friction and because slower mechanical dynamics predominantly influence the release time.
@@
These simulations help establish important design tradeoffs for EA clutches. Because the electroadhesive force is proportional to substrate area, long, narrow EA clutches can achieve much faster engagement and release times while achieving the same force capacity. Diller et al. [6] discovered a similar principle experimentally, when they found that cutting slits into a large EA pad significantly improved release time without harming shear force capacity, and we expect future timing-focused clutch designs will further explore the design space. Improved dynamics also result in faster recovery after slip, opening new uses for EA clutches in applications requiring robustness. The high voltage drive circuit’s rise time also makes a big difference in an EA clutch’s engagement time, which can motivate important decisions over the choice of high voltage amplifier and output stage needed for a project.
VI Conclusions and Future Work
@@
In this paper, we developed a dynamics model for EA-ICE clutches loaded in shear against metallic substrates. As a part of this model, we presented a simple calibration procedure to estimate the interfacial contact mechanics between the dielectric and substrate, bypassing the time and machine cost associated with traditional surface profilometry. We also integrated a dielectric polarization model, proposing that dielectrics can be charged and discharged within microseconds rather than the seconds to minutes commonly assumed in prior modeling work. We then experimentally measured EA shear force capacity, engagement time, and release time over a variety of device parameters and compared our experimental trends to our model’s predictions with good conformity.
@@
We conclude that the drive frequency, voltage, and substrate width to length aspect ratio significantly impact EA engagement and release times. While previous clutches in the literature have focused on square or circular patterns at low drive frequencies (DC or $\leq$10 Hz AC), our simulations and experimental results show that longer, narrower clutches at higher drive frequencies can provide a better trade-off between shear force capacity and actuation time. In our setup, 1-2 kHz drive frequencies achieved the highest shear force capacities, lowest release times, and best slip recoveries, but we expect this optimal range will vary for each EA clutch design.
@@
Our current work has some drawbacks, most notably in its simplified contact mechanics model and handling of parasitics. Future work should investigate more advanced contact models, integrating the substrate’s surface energy and work function to understand what happens at the transition between frictional and peeling fracture mechanics and how the stochastic interactions between local asperities translates to the global shear force. Our model also assumed that the substrate is a metal, in which charges can be induced much faster than in an insulating substrate. Future work should investigate faster and more precise tools to measure the engagement and release times across a variety of substrates.
@@
We believe that improving the dynamics of electroadhesion opens new potential for high-bandwidth soft robots and haptic user interfaces. EA clutches have already received attention among switchable adhesives for their low power, low cost, low mass, and scalable manufacturing processes. However, while micro-electromechanical systems have leveraged these advantages to create kHz-driven precision electrostatic actuators for centimeter-scale rockets and grippers [46, 47], existing soft actuators generally operate under orders of magnitude less accuracy and speed [13]. Similarly, existing EA integrations into haptic user interfaces have historically faced limited bandwidth, with -3 dB frequencies on the order of tens to hundreds of Hertz [23, 10]. By understanding more about how different geometric and operating parameters affect electroadhesion’s response speed, we believe it can become a versatile, space- and power-efficient actuator for future robots.
@@ Acknowledgments
The authors thank Wing-Sum A. Law, Savannah Cofer, Danyang Fan, and Olivia Tomassetti for providing feedback on this paper.

### v2

@@ 4.2 Release Time Results and Discussion
Beyond the parameters varied in Fig. 8, we believe our fast release times are helped by our use of a metallic substrate, which can quickly dissipate induced charge buildup, and our use of a fast push-pull drive circuit output stage, which can quickly sink charge away from the electrodes. Both of these have been used to good effect before for square EA clutches driven by low-frequency drive signals [13, 19, 26, 30], and we believe our exploration of different geometric form factors and high voltage drive signals enable the millisecond-scale release times seen in Fig. 8. We also tested the extent to which our fitted multiplier $\lambda_{ea}^{fit}$ affected our release time simulation results in Supplementary Materials, Sec. S8, and we conclude it had a relatively minor effect for most testing conditions.
4.3 Engagement Time Testing and Simulation Methods
To measure the engagement time, we adapted a method from Diller et al. [7], shown in Fig. 5(d). First, we filtered load cell readings through a zero-phase 250 Hz low-pass Butterworth filter, which was chosen to be similar to prior work [6, 7, 68, 71]. We then linearly interpolated the load cell readings after the voltage is enabled down to the average force over the 10 ms before the voltage is enabled. To minimize the impact of noise, we found linear interpolations for every data interval starting from 0.5 ms after the voltage is enabled and ending between 2-20 ms after the voltage is enabled. After filtering to only fits with $R^{2}>0.8$, we chose the smallest positive interpolated intersection to be the engagement time. For some tests, there would be no positive values due to noise; in these cases, we rounded the run’s measured engagement time up to 0.
4.4 Engagement Time Results and Discussion
Fig. 9 shows the engagement times of our EA-ICE clutches across the same 311 test runs as in Sec. 4.2, showing in general that the EA clutch can engage within several tens of microseconds. This matches our hypothesis when modeling that EA actuators can be designed to engage about as quickly as it takes to polarize their dielectric (for comparison, our power supply had a 10% - 90% rise time of 8.3 µs, and our dielectric had a polarization relaxation time of $\tau=2.82$ µs). Across all 311 test runs, the average engagement time was 4.78 µs ($\sigma$ = 7.07 µs), and 89.7% (279/311) of our measured engagement times were $<$15 µs. We observed significant run-to-run variance in our engagement times, which is not uncommon in the literature given the stochastic nature of the interaction between dielectric asperities and the substrate [7, 9]. For this reason, we focus less on the specific values of our measured engagement times and more on their general trends and order of magnitude.
Comparing results for our $1$ 3.33 cm2 EA-ICE clutches to prior literature, Hinchet and Shea [20] achieved engagement times of 5 ms for a 3 cm2 parallel plate EA clutch, while Zhang et al. [9] achieved an engagement time of 6.7 ms for an 0.96 cm2 EA-ICE clutch adhering to a brass metal substrate. In 1957, Fitch [21] achieved an engagement time of 150 µs for a 38 cm2 parallel plate rotary clutch using silicone oil as the dielectric and very precisely machined electrodes, which we believe helped reduce their interfacial surface roughness. Sub-millimeter scale EA devices like xerographic toner drums [35] and MEMS gap closing actuators [36, 69] also often have microsecond-scale engagement times because, despite having masses an order of magnitude less than that of our system, their initial air gaps are often proportionally much larger than ours due to their fabrication tolerances.
While our engagement times are, on average, better than prior work on centimeter-scale EA clutches, we do not consider them particularly surprising. While the EA release time must account for mechanical damping and load cell dynamics, during engagement the measured shear friction should, theoretically, increase as soon as induced charges accumulate in the substrate, polarizing the dielectric and causing it to reduce the air gap and apply a normal force onto interfacial asperities. Because of our metallic substrate and fast push-pull drive circuit output stage, the electrical dynamics should occur on the order of microseconds, as described in Sec. 2.2. Thus, unlike previous EA dynamics models that predict second-scale time constants due to Maxwell-Wagner interfacial polarization [18, 19, 39], our model correctly matches the order of magnitude of our experimental results. While we are not able to directly apply our model to prior work, given our lack of contact models for their dielectric-substrate interfaces, we explore several hypotheses for why other centimeter-scale EA clutches might see much larger engagement times in Sec. 5.
Although our large run-to-run experimental variance makes direct numerical comparison difficult, our model’s predicted engagement times also generally follow the trends of our experimental data. In Fig. 9(a), linearly interpolating the experimental data gives a similar slope -0.0013 µs V-1 compared to our model’s predicted slope -0.0012 µs V-1. Both Diller et al. [7] and Zhang et al. [9] similarly observed that higher voltages corresponded to lower engagement times. In Fig. 9(b), linearly interpolating the experimental data gives a slope 1.15 µs mm-1 compared to our model’s predicted slope 0.98 µs mm-1. Notably, although the EA normal force is linearly proportional to the substrate width, we see a positive correlation between engagement time and substrate width instead of a negative one because the damping force scales as the cube of substrate width. In Fig. 9(c), as predicted by our model, we did not observe a consistent trend between engagement time and drive frequency because our measured engagement times were less than the period of any of our tested drive signals. In Fig. 9(d), our model predicted a small increase in engagement time for zero preload normal force, but it quickly steadied for higher preload forces once the initial average air gap became small enough that electrical dynamics, rather than mechanical dynamics, dominated our simulation of engagement time. We observed no consistent experimental trend between our measured engagement times and applied preload forces.
5 Simulating Parameter Sweeps for Engagement and Release Time
@@
Fig. 10(a) shows our model’s predicted engagement time over different drive voltages, substrate widths and lengths, dielectric thicknesses, and drive circuit rise times. As shown, the substrate width, drive circuit rise time, and voltage all have very large effects on engagement time, while the dielectric thickness and substrate length have much smaller effects. The substrate’s width affects engagement dynamics much more than its length does because the damping force increases as the cube of width but linearly with length. We also swept a variety of other parameters that did not have a large impact on the predicted engagement time, largely because our actuators engage so quickly that electrical and mechanical dynamics all change on a similar time scale. Our model predicted that the dielectric constant $\kappa$ would have minimal effect on engagement time, because when the voltage first rises (i.e., when $T_{air}$ is the largest) the $\kappa^{2}$ terms in the numerator and denominator of Eq. 4 roughly cancel out. The preload force also had little effect in our predicted and experimental results, which is promising for future work because we found a small preload significantly reduces wrinkling in the dielectric film.
@@
Figure 10: Our model’s predicted (a) engagement times and (b) release times while varying voltage together with (i) substrate width, (ii) dielectric thickness, (iii) substrate length, and (iv) the 10% - 90% rise time or 90% - 10% fall time of our high voltage drive circuit. The black dotted lines indicate the values tested in this paper. Unless otherwise specified, the substrate width tested was 2 mm, the drive signal was a 300 V, 1 kHz bipolar square wave, $F_{preload}=$ 0.125 N, and $F_{shear}(t_{r})/F_{shear,max}$ = 0.8.
@@
Fig. 10(b) shows our model’s predicted release time over the same design parameters. The substrate width matters the most for release time, followed by voltage and dielectric thickness. For larger substrate widths, after the rate of increase in our contact force $F_{k}(T_{air})$ at small air gaps outpaces that of our EA force $F_{ea}(T_{air})$, larger substrate widths are predicted by our model to see lower release times. The dielectric thickness has an interesting local minima in its graph as the dielectric’s increasing mass compensates for the loss of EA force capacity when determining the initial air gap before release. Diller et al. [7] similarly observed higher release times at larger substrate widths and lengths, lower voltages, and in an intermediate dielectric thickness range for parallel plate EA clutches, but to our knowledge this is the first time these trends have been replicated by a physics-based model. The high voltage drive circuitry’s fall time only has a small positive correlation with release time, as it slightly delays when the load cell can overcome static friction and because slower mechanical dynamics predominantly influence the release time. Although not shown in Fig. 10(b), because the electroadhesive force is proportional to the square of both voltage and the dielectric constant $\kappa$, the model similarly predicts that the dielectric constant has a negative correlation with release time. This prediction matches experimental results by Cao et al. [29].
@@
These simulations help establish important design tradeoffs for EA clutches. Because the electroadhesive force is proportional to substrate area, long, narrow EA clutches can achieve much faster engagement and release times while achieving the same force capacity. Diller et al. [7] discovered a similar principle experimentally, when they found that cutting slits into a large EA pad significantly improved release time without harming shear force capacity, and we expect future timing-focused clutch designs will further explore the design space. Improved dynamics also result in faster recovery after slip, opening new uses for EA clutches in applications requiring robustness. The high voltage drive circuit’s rise time also makes a large difference in an EA clutch’s engagement time, which can motivate important decisions over the choice of high voltage amplifier and output stage needed for a project.
6 Conclusions and Future Work
@@
In this paper, we developed a dynamics model for EA-ICE clutches loaded in shear against metallic substrates. As a part of this model, we presented a simple calibration procedure to estimate the interfacial contact mechanics between the dielectric and substrate, bypassing the time and machine cost associated with traditional surface profilometry. We also integrated a Cole-Cole dielectric polarization model, proposing that dielectrics can be charged and discharged within microseconds rather than the seconds to minutes commonly assumed in prior modeling work. We then experimentally measured EA shear force capacity, release time, and engagement time over a variety of device parameters and compared our experimental trends to our model’s predictions with good conformity.
@@
We conclude that the drive frequency, voltage, and substrate width to length aspect ratio significantly impact EA release times. While previous clutches in the literature have focused on square or circular patterns at low drive frequencies (DC or $\leq$10 Hz AC), our simulations and experimental results show that longer, narrower clutches at higher drive frequencies can provide a better trade-off between shear force capacity and release time. In our setup, 1-2 kHz drive frequencies achieved the highest shear force capacities, lowest release times, and best slip recoveries, but we expect this optimal range will vary for each EA clutch design. We also find, both experimentally and in simulation, that our use of a metal substrate and fast push-pull high voltage output stage significantly improved EA engagement time, allowing our EA clutches to engage about as fast as their dielectric can polarize across a wide range of geometric and operating parameters.
@@
Our current work has some drawbacks, most notably in its simplified contact mechanics model and handling of parasitics. Future work should investigate more advanced contact models, integrating the substrate’s surface energy and work function to understand what happens at the transition between frictional and peeling fracture mechanics, how the stochastic interactions between local asperities translates to the global shear force, and how surface roughness factors into the electroadhesive force as asperities compress into the soft dielectric. The ultimate ideal is a model with no fitted multiplier $\lambda_{ea}^{fit}(V)$. Our model also assumed that the substrate is a metal, in which charges can be induced much faster than in an insulating substrate. Future work should investigate faster and more precise tools to measure the engagement and release times across a variety of substrates. Optical interferometry [42] or digital image correlation [72], for example, could allow for higher sample rate, more spatially nuanced measurements of how asperities on the two surfaces adhere and release over time. In lieu of perfectly accurate physics-based models, an alternative approach could be a completely data-driven model for electroadhesion, similar to the one explored by Diller et al. [7]. In addition, although we originally used lower preload forces to avoid plastically deforming the dielectric film, our finding that lower preload forces tend to improve release times opens interesting translational questions into how to best integrate EA clutches in the field. Prior studies have explored a variety of elastic preload mechanisms for different EA clutch geometries [7, 12, 13, 73], and additional topology optimization should explore optimal mechanisms that can minimize wear and release time while still ensuring good surface conformity.
@@
We believe that improving the dynamics of electroadhesion opens new potential for high-bandwidth applications, such as soft robots and haptic user interfaces. EA clutches have already received attention among switchable adhesives for their low power, low cost, low mass, and scalable manufacturing processes. However, while micro-electromechanical systems have leveraged these advantages to create kHz-driven precision electrostatic actuators for centimeter-scale rockets and grippers [36, 74], existing soft actuators generally operate under orders of magnitude less accuracy and speed [14]. Similarly, existing EA integrations into haptic user interfaces have historically faced limited bandwidth, with -3 dB frequencies on the order of tens to hundreds of Hertz [11, 41]. By understanding more about how different geometric and operating parameters affect electroadhesion’s response speed, we believe it can become a versatile, space- and power-efficient actuator for future robotic systems.
@@ Acknowledgments
This work is supported by the National Science Foundation award grant no. 2142782. The authors thank Wing-Sum Adrienne Law, Savannah Cofer, Danyang Fan, and Olivia Tomassetti for providing feedback on this paper. The authors also thank Hojung Choi and Kai Jun Chen for their insightful discussions on capacitive sensing and fracture mechanics, respectively.